# Mirror-3DGS: Incorporating Mirror Reflections into 3D Gaussian Splatting

### v1

@@ 3.4 Two-Stage Training Strategy
where $\boldsymbol{p}_{i}$ is the coordinate of the Gaussian to be optimized, $N_{f}$ is the number of filtered mirror Gaussians.
Therefore, the overall loss for the first stage is:
@@
where $\lambda_{mask},\lambda_{depth}$ are the weight of mask loss and depth loss.
Stage 2: The objective of the second stage is to refine the rendering quality of the fused images, during which the mirror equation is fixed. As described in Sec. 3.3, we fuse images from the current viewpoint and the virtual mirrored viewpoint, supervising with the complete, unmasked GT image $\mathbf{C}_{gt}$:
$\mathcal{L}_{rgb}^{2}=(1-\gamma)\mathcal{L}_{1}(\mathbf{C}_{fuse},\mathbf{C}_{gt})+\gamma\mathcal{L}_{D-SSIM}(\mathbf{C}_{fuse},\mathbf{C}_{gt}),$ (15)
@@
$\mathcal{L}_{\text{stage}}^{2}=\lambda_{mask}\mathcal{L}_{mask}+\mathcal{L}_{rgb}^{2}+\lambda_{depth}\mathcal{L}_{depth}.$ (16)
Table 1: Quantitative comparison on the complete test set. The first three are NeRF-based methods and Evoluted-3DGS is our baseline described in Sec. 3.2, best results bolded and second best underlined. Methods marked with an asterisk ‘*’ are evaluated solely on the mirror regions.
@@
Mirror-3DGS* (Ours) 32.40 - - - 22.51 - - -
Table 2: Quantitative comparison on the challenging test set with all images containing mirrors. The first three are NeRF-based methods and Evoluted-3DGS is our baseline described in Sec. 3.2, best results bolded and second best underlined. Methods marked with an asterisk ‘*’ are evaluated solely on the mirror regions.
Method Synthetic Real
PSNR↑ SSIM↑ LPIPS↓ FPS↑ PSNR↑ SSIM↑ LPIPS↓ FPS↑
InstantNGP [28] 14.26 0.44 0.65 9.22 10.54 0.20 0.72 13.00
DVGO [33] 17.65 0.63 0.49 4.34 22.28 0.68 0.32 4.41
Mirror-NeRF [49] 27.33 0.94 0.05 0.55 25.25 0.86 0.06 0.55
3DGS [18] 22.15 0.84 0.17 275.73 22.13 0.72 0.24 208.23
Evoluted-3DGS 21.82 0.82 0.09 195.36 21.59 0.75 0.11 98.49
Mirror-3DGS (Ours) 24.73 0.85 0.06 120.61 23.01 0.77 0.09 95.22
3DGS* [18] 16.42 - - - 21.27 - - -
Mirror-NeRF* [49] 22.14 - - - 22.18 - - -
Mirror-3DGS* (Ours) 21.60 - - - 22.94 - - -
4 Experiments
4.1 Experimental Setup
Dataset. We utilized the publicly released dataset by Mirror-NeRF [49], encompassing common scenarios with prevalent mirrors. Specifically, the dataset includes three synthetic scenes (living room, office, and washroom) and three real-world captured scenes (discussion room, lounge, and market). The number of training views per scene ranges from 200 to 300, each accompanied by a mirror mask. Each scene is equipped with a complete test set featuring a mix of views with and without mirrors, alongside a challenging test set exclusively composed of mirror-containing images.
Implementation. Our Mirror-3DGS was developed based on the codebase of 3DGS [18]. Training and testing are conducted on 4×4$\times$4 × downsampled resolution (i.e., $480\times 360$) for all three real-world scenes, and 2×2$\times$2 × downsampled resolution (i.e., $400\times 400$ for living room and $400\times 300$ for the others) for synthetic scenes. The training process is divided into two stages: in the first stage, we train for 5000 steps, followed by 65000 steps in the second stage. The loss weights $\lambda_{mask}$ and $\lambda_{depth}$ are set to 1 and 0.1 in all experiments, respectively. Our code will be made publicly available for reproducible research.
@@ 4.2 Experimental Results
Quantitative Comparisons. We conduct quantitative comparisons between our method and several competitive counterparts, including InstantNGP [28], DVGO [33], the state-of-the-art Mirror-NeRF [49], vanilla 3DGS [18] and our baseline Evoluted-3DGS. The comparison results across the complete and the challenging test sets are detailed in Tab. 1 and Tab. 2, respectively.
@@
Among these competitors, Mirror-NeRF consistently achieves the highest PSNR, SSIM, and LPIPS scores on both synthetic and real scenes. However, its rendering speed, i.e., FPS, is significantly lower due to the inherent inefficiencies of NeRF’s volumetric rendering mechanism. Even the fastest NeRF-based method among these, InstantNGP, is orders of magnitude slower compared to 3DGS and our approach in real scenes. As such, these NeRF-based methods are challenging to apply in scenarios requiring real-time rendering.
@@
Our method yields performance slightly better than 3DGS within the complete test set, which includes a minimal number of views with mirrors, yet it significantly surpasses vanilla 3DGS in the challenging test set, where mirror views are predominant. Specifically, our Mirror-3DGS demonstrates an average PSNR improvement of 1.58 and 0.8 dB over vanilla 3DGS on the challenging test set of synthetic and real scenes, respectively. Remarkably, our method even surpasses Mirror-NeRF in washroom scene, achieving a PSNR gain of 4.88 dB. Additionally, our SSIM and LPIPS metrics far exceed those of vanilla 3DGS, indicating that our synthesized views are significantly more realistic. These improvements are mainly attributed to our method’s specialized handling of mirrors and their reflections in the scene, which inevitably introduces some additional computation overhead. Therefore, while our rendering speed is slightly lower than that of vanilla 3DGS, it is significantly faster than NeRF-based methods, making it suitable for real-time rendering applications.
@@
It is important to highlight that our Mirror-3DGS outperforms both Mirror-NeRF and vanilla 3DGS in terms of the PSNR metric when focusing solely on the mirror regions, as shown in the last three rows of Tab. 1 and Tab. 2, demonstrating our method’s enhanced capability in handling mirror reflections.
Figure 4: Qualitative comparison of novel view synthesis. Rows 1-6 respectively showcase rendering results of living room, office, washroom, discussion room, lounge and market scenes, with magnified displays of mirror regions in the top-right and predicted mirror mask in the bottom-right of the rendered views. The PSNR is computed solely on the mirror region. Mirror-NeRF and Mirror-3DGS can accurately predict the mirror masks and our Mirror-3DGS achieves the best performance on mirror regions.
Figure 5: Evolution of rendering results over training time on the washroom scene. Our Mirror-3DGS rapidly learns to accurately render contents in mirrors, while Mirror-NeRF takes more than five hours to reach similar levels of detail.
Figure 6: Convergence speed comparison on the washroom scene. Our method achieves optimal rendering results in approximately 15 minutes, significantly outpacing Mirror-NeRF, which requires over 8 hours to reach a comparable level of performance.
Qualitative Comparisons. The qualitative analysis presented in Fig. 4, focusing on the detailed rendering of mirrors, contrasts our Mirror-3DGS method with three competitive alternatives across six diverse scenes.
The grid-based DVGO method often results in blurry images of mirror content across all scenes, due to limitations in grid resolution and inaccuracies in physical modeling that cause distortions and artifacts on mirror surfaces. The 3DGS method also struggles with accurately rendering mirrors. Its main problem is the incorrect placement of Gaussians for content within the mirrors, positioned behind the mirror, which distorts the scene’s depth information and leads to inaccurate renderings. Mirror-NeRF, utilizing ray tracing to render content within mirrors, more accurately captures physical imaging principles, achieving the highest overall rendering quality. However, its method for implicitly modeling the probabilities of mirror reflections and normals can sometimes lead to inaccuracies in depicting the reflective content of mirrors. Furthermore, integrating ray tracing with the 3DGS is challenging.
@@
Our Mirror-3DGS, by incorporating the principles of plane mirror imaging and accurately modeling spatial geometric relationships, enables high-fidelity rendering of mirror contents. Across all six scenes, our Mirror-3DGS achieves rendering quality comparable to Mirror-NeRF, demonstrating its effectiveness in handling complex mirror reflection.
Training Time and Convergence. A significant advantage of our method is the real-time rendering capability derived from 3DGS. In parallel, our method also boasts a substantial advantage in training efficiency compared to NeRF-based counterparts, particularly against the current state-of-the-art method, Mirror-NeRF. Fig. 5 and Fig. 6 showcase the time-evolving rendering effects on the washroom scene. It is evident that our method requires only 1 minute of training to achieve satisfactory rendering of the contents within the mirror, whereas Mirror-NeRF starts to render the mirror’s contents only after 5 hours of training. Ultimately, our method converges to its best performance after 15 minutes of training, while Mirror-NeRF takes 8 hours to converge. Beyond the advantage in convergence speed, our method closely matches the visual quality of Mirror-NeRF and even surpasses it in some scenes.
4.3 Ablation Study
Our Mirror-3DGS method incorporates several key designs: 1) $\mathcal{L}_{mask}$, 2) $\mathcal{L}_{depth}$, and 3) $\mathcal{L}_{plane}$. Through ablation results displayed in Tab. 3 and Tab. 4, we confirm their significance to our approach.
Firstly, the design of $\mathcal{L}_{mask}$ is crucial for effectively learning mirror properties, enabling precise identification of mirror regions during rendering. Removing $\mathcal{L}_{mask}$ leads to the significant performance drop, attributed to incorrect estimation of mirror areas, which not only affects the rendering of content within mirrors but also results in the erroneous processing of non-mirror regions as if they were mirrors, thereby producing incorrect rendering results.
Secondly, $\mathcal{L}_{depth}$ ensures consistency between the rendered depth in the learning process and the depth extracted by the depth model. Given that the pre-trained depth model is relatively accurate in handling the depth of mirror regions, it effectively prevents the incorrect generation of Gaussians for content within mirrors as if they were physically present. The absence of this supervision noticeably reduces our method’s performance.
Table 3: Ablation results on the challenging test set of real market scene. The values in red parentheses indicate metric changes relative to our full method.
PSNR↑ SSIM↑ LPIPS↓
w/o $\mathcal{L}_{mask}$ 25.68 (↓1.50) 0.83 (↓0.01) 0.05 (↑0.01)
w/o $\mathcal{L}_{depth}$ 25.39 (↓1.79) 0.82 (↓0.02) 0.05 (↑0.01)
w/o $\mathcal{L}_{plane}$ 26.84 (↓0.34) 0.84 (↓0.00) 0.04 (↑0.00)
Our full method 27.18 (↓0.00) 0.84 (↓0.00) 0.04 (↑0.00)
Table 4: Ablation results on the challenging test set of synthetic washroom scene. The values in red parentheses indicate metric changes relative to our full method.
PSNR↑ SSIM↑ LPIPS↓
w/o $\mathcal{L}_{mask}$ 17.27 (↓3.03) 0.69 (↓0.06) 0.13 (↑0.04)
w/o $\mathcal{L}_{depth}$ 18.61 (↓1.69) 0.72 (↓0.03) 0.11 (↑0.02)
w/o $\mathcal{L}_{plane}$ 16.29 (↓4.01) 0.67 (↓0.08) 0.15 (↑0.06)
Our full method 20.30 (↓0.00) 0.75 (↓0.00) 0.09 (↑0.00)
Finally, $\mathcal{L}_{plane}$ aids in learning more accurate mirror planes. If the learned mirror plane is rough, applying the transformation in Eq. (8) leads to incorrect imaging results in the mirror. We observed that this loss has a minor impact on performance on real scenes, while its influence is more pronounced in synthetic scenes.
5 Conclusion
@@
In this paper, we introduced Mirror-3DGS, a novel rendering framework that addresses the challenge of accurately capturing and rendering physical reflections in mirror-containing scenes. Our Mirror-3DGS employs a two-stage training process that first filters mirror Gaussians and then uses this information to accurately estimate the mirror plane and derive mirrored camera parameters. The effectiveness of Mirror-3DGS in delivering high-quality renderings of scenes with mirrors is demonstrated by experiments on three synthetic and three real scenes, closely matching the state-of-the-art Mirror-NeRF method, while achieving substantial improvements in training and rendering speed. However, our Mirror-3DGS still faces limitations, notably in the accuracy of mirror estimations and its dependence on dataset-provided mirror masks. Currently, our method is tailored for single flat mirrors, leaving the adaptation to complex shapes and multiple mirrors as an objective for future exploration.
Supplementary Material
This supplementary material provides additional details of our proposed method, including formula derivation in Sec. 0.A, and method details in Sec. 0.B.
Appendix 0.A Mirror Transformation Matrix
The relationship between the viewpoints $\mathbf{P_{o}}$ and $\mathbf{P_{m}}$ is derived here, corresponding to main paper Eq. (8). As described in main paper Eq. (7), an infinite plane can be parameterized as:
$\boldsymbol{n}_{\pi}^{\top}\boldsymbol{p}+d=0.$ (17)
where the unit vector $\boldsymbol{n}_{\pi}$$=$$(a,b,c)^{\top}$ represents the normal of the plane and $\boldsymbol{p}=(X,Y,Z)^{\top}$ is a point upon the plane. The Eq. (17) can also be written as:
$aX+bY+cZ+d=0.$ (18)
Let $(x,y,z)^{\top}$ be a point outside the mirror plane, then the projection of this point onto the plane is given by:
$\begin{bmatrix}x_{p}\\ y_{p}\\ z_{p}\\ \end{bmatrix}=\begin{bmatrix}x\\ y\\ z\\ \end{bmatrix}-\lambda{\begin{bmatrix}a\\ b\\ c\\ \end{bmatrix}},$ (19)
where $\lambda=ax+by+cz+d$.
According to the principle of mirror reflection, the coordinates of the mirror-symmetrical point $(x_{r},y_{r},z_{r})^{\top}$ can be calculated as:
$\begin{bmatrix}x_{r}\\ y_{r}\\ z_{r}\\ \end{bmatrix}=\begin{bmatrix}x_{p}\\ y_{p}\\ z_{p}\\ \end{bmatrix}-\lambda{\begin{bmatrix}a\\ b\\ c\\ \end{bmatrix}}=\begin{bmatrix}x\\ y\\ z\\ \end{bmatrix}-2\lambda{\begin{bmatrix}a\\ b\\ c\\ \end{bmatrix}},$ (20)
The corresponding homogeneous coordinate representation is:
$\begin{bmatrix}x_{r}\\ y_{r}\\ z_{r}\\ 1\end{bmatrix}=\begin{bmatrix}x\\ y\\ z\\ 1\\ \end{bmatrix}-2\lambda{\begin{bmatrix}a\\ b\\ c\\ 0\\ \end{bmatrix}}=\begin{bmatrix}1-a^{2}&-2ab&-2ac&-2ad\\ -2ab&1-2b^{2}&-2bc&-2bd\\ -2ac&-2bc&1-2c^{2}&-2cd\\ 0&0&0&1\end{bmatrix}\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix},$ (21)
We define $\mathbf{T_{m}}$ as:
$\mathbf{T_{m}}=\begin{bmatrix}1-a^{2}&-2ab&-2ac&-2ad\\ -2ab&1-2b^{2}&-2bc&-2bd\\ -2ac&-2bc&1-2c^{2}&-2cd\\ 0&0&0&1\end{bmatrix}.$ (22)
corresponding to main paper Eq. (9).
Similarly, the relation between $(x,y,z)^{\top}$ and $(x_{r},y_{r},z_{r})^{\top}$ can be written as:
$\mathbf{Q_{o}}{\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix}}=\mathbf{Q_{m}}{\begin{bmatrix}x_{r}\\ y_{r}\\ z_{r}\\ 1\end{bmatrix}},$ (23)
where $\mathbf{Q_{o}}$ and $\mathbf{Q_{m}}$ are original and mirrored viewpoint respectively, representing the transformation from world coordinates to camera coordinates. Then,
$\mathbf{Q_{o}}{\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix}}=\mathbf{Q_{m}}\mathbf{T_{m}}{\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix}}.$ (24)
Since $(x,y,z)^{\top}$ represents any arbitrary point, we can deduce that:
$\mathbf{Q_{o}}=\mathbf{Q_{m}}\mathbf{T_{m}}.$ (25)
Finally, the relation between their inverse matrices $\mathbf{P_{o}}$ and $\mathbf{P_{m}}$, as illustrated in Eq. (8) of our main paper, can be written as:
$\mathbf{P_{m}}=\mathbf{T_{m}}\mathbf{P_{o}}.$ (26)
Appendix 0.B Method Details
0.B.1 The detail of rendering
In the Section 3.3 of the main paper, we elaborate on the virtual mirrored viewpoint construction and image fusion. Here, we represent the detailed algorithmic procedure as pseudocode, which is displayed in Algorithm 1.
Algorithm 1 Reflective Rendering
  Input: original viewpoint $\mathbf{P_{o}}\in\mathbb{R}^{4\times 4}$, gaussians $\mathbf{G}\in\mathbb{R}^{N\times C}$, threshold $\tau$.
  Output: the fused view image $\mathbf{C}_{fuse}\in\mathbb{R}^{H\times W\times 3}$.
  # $\mathbf{C}_{\mathbf{P_{o}}},\mathbf{D}_{pred},\mathbf{M}$ are rendered image, depth map, and mirror mask, respectively
  $\mathbf{C}_{\mathbf{P_{o}}},\mathbf{D}_{pred},\mathbf{M}\leftarrow Render(\mathbf{P_{o}},\mathbf{G})$
  $\mathbf{G}_{\mathbf{mirror}}\leftarrow MirrorFilter(\mathbf{G},\tau)$ Gaussian points represent mirror plane
  $a,b,c,d\leftarrow RANSAC(\mathbf{G}_{\mathbf{mirror}})$ mirror plane parameters described in Eq. (18)
  $\mathbf{C}_{\mathbf{P_{m}}}\leftarrow Render(\mathbf{P_{m}},\mathbf{G})$ Render image of mirrored viewpoint
  $\mathbf{C}_{fuse}=\mathbf{C}_{\mathbf{P_{o}}}\odot{(\mathbbm{1}-\mathbf{M}})+\mathbf{C}_{\mathbf{P_{m}}}\odot\mathbf{M}$ Image fusion, Eq. (10) in the main paper
  return  $\mathbf{C}_{fuse}$
0.B.2 Mirror Plane Estimation
To learn the accurate parameters of the mirror plane, we introduce $\mathcal{L}_{plane}$ to constrain the Gaussians constituting the mirror surface.
In main text, we employ the RANSAC[11] method for plane selection and use the average distance from points to the mirror surface as $\mathcal{L}_{plane}$. However, for scenes where the mirror is distant from the camera center, significant distortion occurs, which RANSAC cannot process well.
In order to handle the distortion, we develop another way to choose our plane. By selecting the median normal vector and median coordinates of the mirror points as the normal of and a point upon the plane, we can determine the coordinates of the mirror. For above reasons, we need to constrain the normal vectors.
According to [31], we consider the short-axis direction of the Gaussian as its normal vector. Constraints to the normal vectors are appended to ensure them to be parallel. Specifically, we randomly select three mirror Gaussians and require that their normal vectors, denoted as $\mathbf{n_{a},n_{b},n_{c}}$, are parallel to each other:
$\mathcal{L}_{plane}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}|\mathbf{n_{a}}\times\mathbf{n_{b}}\cdot\mathbf{n_{c}}|,$ (27)
where $N_{p}$ denotes the number of the 3-points sets randomly selected from mirror plane, $\times$ represents the cross product, and $\cdot$ represents the dot product.
The rest of the approach remains the same as mentioned in the main text. We recommend using the RANSAC to determine the plane in scene discussion room, lounge and market, while using the approch mentioned above to determine the rest scene.
0.B.3 Activation Function
We use the SINE activation function instead of the activation function Sigmoid in 3DGS. The expression of SINE is as follows:
${SINE(x)}=\frac{sin(x)+1}{2}.$ (28)
We apply SINE and Sigmoid activation function to normalize the opacity and mirror properties respectively, both of which are bounded within the range of [0, 1], while keeping all other experimental variables constant.
Figure 7: Illustration of Convergence Speed and Storage Size for the Washroom Scene. The radius of points on the line indicates the storage size of the scene. The SINE activation function accelerates convergence and significantly reduces the number of Gaussians, i.e., lowering storage cost.
As depicted in Fig. 7, the empirical findings indicate that SINE activation function significantly enhances the model’s convergence rate and diminishes its complexity. This improvement is attributed to the SINE function’s circumvention of intricate exponential computations. Moreover, in contrast to Sigmoid function, SINE function mitigates the vanishing gradient issue and moderates the localization gradient, thereby curtailing the superfluous densification process.

### v2

@@ II-D Two-Stage Training Strategy
where $\boldsymbol{p}_{i}$ is the coordinate of the Gaussian to be optimized, $N_{f}$ is the number of filtered mirror Gaussians. The overall loss for the first stage is:
@@
where $\lambda_{mask},\lambda_{depth}$ are the weight of mask loss and depth loss, respectively.
Stage 2: The objective of the second stage is to refine the rendering quality of the fused images, during which the mirror equation is fixed. As described in Sec. II-C, we fuse images from the current viewpoint and the virtual mirrored viewpoint, supervising with the complete, unmasked GT image $\mathbf{C}_{gt}$:
$\mathcal{L}_{rgb}^{2}=(1\!-\!\gamma)\mathcal{L}_{1}(\mathbf{C}_{fuse},\mathbf{C}_{gt})+\gamma\mathcal{L}_{D-SSIM}(\mathbf{C}_{fuse},\mathbf{C}_{gt})$ (14)
@@
$\mathcal{L}_{\text{stage}}^{2}=\lambda_{mask}\mathcal{L}_{mask}+\mathcal{L}_{rgb}^{2}.$ (15)
Figure 4: Qualitative comparison of novel view synthesis. Two raws display washroom and discussion room scenes. The top-right shows magnified mirror regions, and the bottom-right shows predicted mirror masks. PSNR is calculated only for the mirror regions. Both Mirror-NeRF and Mirror-3DGS accurately predict mirror masks, with Mirror-3DGS achieving the best performance.
TABLE I: Quantitative comparison on the Mirror-NeRF dataset. Best results are bolded and second best are underlined. Methods marked with ‘*’ are evaluated only on mirror regions.
@@
Mirror-3DGS* (Ours) 32.40 - - - 24.53 - - -
III Experiments
III-A Experimental Setup
Dataset. We utilize the publicly released dataset by Mirror-NeRF [7], which includes common scenarios with prevalent mirrors. The dataset comprises three synthetic scenes (living room, office, and washroom) and three real-world scenes (discussion room, lounge, and market). Each scene has 200 to 300 training views with accompanying mirror masks.
Implementation. Training and testing are done at $480\times 360$ resolution for real scenes and $400\times 400$ for synthetic scenes. The training process is divided into two stages: we train 5000 steps in stage 1, followed by 65000 steps in stage 2. The loss weights $\lambda_{mask}$ and $\lambda_{depth}$ are set to 1 and 0.1, respectively.
@@ III-B Experimental Results
Quantitative Comparisons. We conduct quantitative comparisons between our method and several competitive counterparts, and the comparison results are detailed in Tab. I.
@@
Among these competitors, Mirror-NeRF achieves high PSNR, SSIM, and LPIPS scores in synthetic and real scenes. However, its rendering speed is significantly lower due to NeRF’s volumetric rendering inefficiencies. Even the fastest NeRF-based method, InstantNGP, is much slower than 3DGS and our approach in real scenes. Therefore, NeRF-based methods are challenging for real-time rendering scenarios.
@@
Our method yields performance better than vanilla 3DGS. Specifically, our Mirror-3DGS demonstrates an average PSNR improvement of 0.89 and 3.38 dB over vanilla 3DGS on the synthetic and real scenes, respectively. Additionally, our SSIM and LPIPS metrics far exceed those of vanilla 3DGS, indicating that our synthesized views are significantly more realistic. These improvements are mainly attributed to our method’s specialized handling of mirrors and their reflections in the scene, which inevitably introduces some additional computation overhead. Therefore, while our rendering speed is slightly lower than that of vanilla 3DGS, it is significantly faster than NeRF-based methods, making it suitable for real-time rendering applications.
@@
It is important to highlight that our Mirror-3DGS outperforms both Mirror-NeRF and vanilla 3DGS in terms of the PSNR metric when focusing solely on the mirror regions, as shown in the last three rows of Tab. I, demonstrating our method’s enhanced capability in handling mirror reflections.
@@
Qualitative Comparisons. The qualitative analysis presented in Fig. 4, focusing on the detailed rendering of mirrors. Our Mirror-3DGS, by incorporating the principles of plane mirror imaging and accurately modeling spatial geometric relationships, enables high-fidelity rendering of mirror contents. Across all six scenes, our Mirror-3DGS achieves rendering quality comparable to Mirror-NeRF, demonstrating its effectiveness in handling complex mirror reflection.
IV Conclusion
@@
In this paper, we introduce Mirror-3DGS, a novel rendering framework that addresses the challenge of accurately capturing and rendering physical reflections in mirror-containing scenes. Our Mirror-3DGS employs a two-stage training process that first filters mirror Gaussians and then uses this information to accurately estimate the mirror plane and derive mirrored camera parameters. The effectiveness of Mirror-3DGS in delivering high-quality renderings of scenes with mirrors is demonstrated by experiments on synthetic and real scenes.